Table 2.1 contains information on three techniques for producing $15 worth of bar soap. Assume that we specified “$15 worth of bar soap” because soap costs $3 per bar and all three techniques produce 5 bars of soap ($15 = $3 per bar × 5 bars). So you know each technique produces 5 bars of soap.



a. What technique will you want to use if the price of a bar of soap falls to $2.75? Which technique will you use if the price of a bar of soap rises to $4? To $5?



b. How many bars of soap will you want to produce if the price of a bar of soap falls to $2.00?



c. Suppose that the price of soap is again $3 per bar but that the prices of all four resources are now $1 per unit. Which is now the least-profitable technique?



d. If the resource prices return to their original levels (those shown in the table) but a new technique is invented that can produce 3 bars of soap (yes, 3 bars, not 5 bars!) using 1 unit of each of the four resources, will firms prefer the new

1
Expert's answer
2021-11-17T11:05:09-0500

a) If the price of bar soap falls to $2.75, then 5 bars cost; $2.75 *5 = $13.75.

Technique 2 has a production cost lower than $13.75 so that is the one that should be used.


When the price rises to $4 per bar, 5 bars cost; $4 *5 =$20

Any of the techniques can be used, but since technique 2 has the lowest production cost, this is the one that should be used.


If the price rises to $5, technique 2 should still be used. The lowest production costs will yield the largest profit.


b) If the price of a bar of soap is only $2, the lowest production costs are needed.

It takes $13 to make 5 bars worth ; $5 *2 =$10

It takes $26 to make 10 bars worth ; $10 *2 = $20

It takes $39 to make 15 bars worth : $15 *2 = $30

Using this logic, there is no way for sales to meet or beat production costs. Changes would have to be made to production.


c) If the cost of each resource is reduced to $1, the least profitable technique is the one that requires the most units of resources. From the table, both technique 1 and 2 require 7, and technique 3 requires 8.Technique 3 is therefore the least profitable.


d) If the cost of all the resources is $1, the cost to produce 3 bars of soap is;

= 2(1) + 1(1) + 3(1) + 3(1)

= 2 + 1 + 3 + 3

= 9

This is the same cost as techniques 1 and 3 for 5 bars. Technique 2 is still the least expensive option.